OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of three-sevenths of the first number and one-third of the second number is equal to thirty-eight. i got 49 and 51.

OpenStudy (anonymous):

Well Done you are right..

OpenStudy (anonymous):

\[\frac{3}{7} \times 49 + \frac{1}{3} \times 51 \implies 3 \times 7 + 17 \implies 21 + 17 = 38\]